OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = -625 and a2 = 125? Select one: A. -0.2 B. 0.04 C. -0.04 D. 0.2

OpenStudy (anonymous):

@mechanicalanimal_

OpenStudy (anonymous):

@iambatman

OpenStudy (cwrw238):

first find the common ratio which is 125 / -625

OpenStudy (cwrw238):

got that?

OpenStudy (anonymous):

Yea i think so

OpenStudy (cwrw238):

what do you get?

OpenStudy (anonymous):

That the common help you determine how many are between each number

OpenStudy (anonymous):

they want the seventh term

OpenStudy (cwrw238):

the coomon ratio = -0.2 7th term = a1*r^6 = -625 * (-0.2)^6

OpenStudy (cwrw238):

sorry i was wrong - can you work out -625*(-0.2)^6 on ur calculator

OpenStudy (anonymous):

i dont have a calculator

OpenStudy (cwrw238):

there should be one on your computer

OpenStudy (anonymous):

it doesnt let me do ^6

OpenStudy (anonymous):

Is it C?

OpenStudy (cwrw238):

you are probably in the standard mode go into view and press scientific mode

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

Thanks!